pearl divers from Ceylon can go down to a depth of 17 fathoms without breathing gear. convert this depth to meters(1 fathom=6 feet. What is the the weight density for water and how much water pressure is on them at that depth.
Answer #1

1 fathom=6 feet

So, 17 fathoms = 17*6 ft = 102 ft

Now, 1 ft = 0.3048 m

So, 17 fathoms = 102*0.3048 m = 31.1 m (answer)

Density of water = 1000 kg/m^3 (ans)

Pressure = density*g*h

= 1000*9.8*31.1

= 304,679 Pa (ans)